answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Name Three possible outcomes of obesity?

Diabities, blood clot, being made fun of, embarrassment, appearance ect.

How come you do not need to have three actual coins to compute the probabilities for these outcomes?

Because the theory of coin flipping is well understood and so theoretical probabilities can be used.

How many distinct three letter arrangements can be found from the letters in mathematics?


The number of different three letter arrangements that can be done from the

letters in the word "mathematics"

is; 11P3 =

11!/(11-3)! =

990

Keeping score of a set of observations?

you could use a table saying when you saw it on 1 side and where on the other side

What is the probability of rolling two dice that the first dice is an even number?

There are only three even numbers on a "legal" six sided dice (2,4 & 6), which ever dice is designated as "first" whether rolled independently or in unison will have a 1 in 2 chance of being even.

Put another way: legal dice, designate first, 3 even, 3 odd, 50/50.

Bonus Answer: Applies to second dice also.

What is the best way to study probability in math?

It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.

It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.

It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.

It is best to study it systematically and carefully. Make sure you understand the various concepts before moving on to more complicated ideas.

What is the estimd population mean of sample 36 mean weight 3.01 standard diviation 03?

The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.

The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.

The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.

The answer will depend on the population mean of what variable? Height?, length or is it simply weight. If it is weight, the estimated (not estimd) population mean is 3.01 units: the same as the sample mean. The standard deviation (not diviation) is irrelevant.

What is the probability that they will have at least one girl if a couple plans to have 6 children?

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then

Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)

= 1 - Prob(6 boys out of 6 children)

= 1 - 0.526 = 0.9809

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then

Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)

= 1 - Prob(6 boys out of 6 children)

= 1 - 0.526 = 0.9809

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then

Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)

= 1 - Prob(6 boys out of 6 children)

= 1 - 0.526 = 0.9809

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a girl is approx 0.48, then

Prob(at least one girl out of 6 children) = 1 - Prob(no girls out of 6 children)

= 1 - Prob(6 boys out of 6 children)

= 1 - 0.526 = 0.9809

What is the function of the chromatain?

The chromatain have four major functions. They package DNA into a smaller volume to fit in the cell. They strengthen the DNA to allow mitosis, and they prevent damage to DNA. Chromatain control gene expression and DNA replication.

What is the function of the common bileduct?

To drain bile from the liver (and the gall bladder) into the small intestines.

(Bile is a combination of waste products (mostly from decomposing hemoglobin) and solvents designed to help dissolve (digest) fats in your food.)

What are the chances of becoming a freelance writer in California?

Freelancing is a home based job, you can do it from any where and serve customers of any location.

What you need to have is excellent writing skills and analytical skills. If you are looking for working as a freelancer, you can choose Obuhi and get hired as per your skills they will help you to find a good job.

Suppose you roll two ordinary 6 sided dice how may ways could the product of the dice be less then12?

(1,1) = 1 (1,2) = 2 (1,3) = 3 (1,4) = 4 (1,5) = 5 (1,6) = 6

(2,1) = 2 (2,2) = 4 (2,3) = 6 (2,4) = 8 (2,5) = 10 (2,6) = 12

(3,1) = 3 (3,2) = 6 (3,3) = 9 (3,4) = 12 (3,5) = 15 (3,6) = 18

(4,1) = 4 (4,2) = 8 (4,3) = 12 (4,4) = 16 (4,5) = 20 (4,6) = 24

(5,1) = 5 (5,2) = 10 (5,3) = 15 (5,4) = 20 (5,5) = 25 (5,6) = 30

(6,1) = 6 (6,2) = 12 (6,3) = 18 (6,4) = 24 (6,5) = 30 (6,6) = 36

Out of 36 possibilities, 6 + 5 + 3 + 2 + 2 + 1 = 19 fit the requirements.

What do you notice about the distribution of volcaneos?

They tend to occur over pockets of magma and along the edges of tectonic plates

Trending Questions
What are the odds of not rolling a 7 with a pair of 6 sided dice in 23 rolls? Why is probability theory important? Jerry has a total of 23 marbles The marbles are either blue or green He has three more blue marbles than green marbles How many green marbles? What is the approximate distance an electron is located from the nucleus measured by? What is percentile deviation? What are the odds of a number less than 2 when rolling a die? How do you draw a club from a deck of cards? When the population standard deviation is unknown the sampling distribution is equal to what? What is the probability of drawing a king and a diamond from a deck of playing cards? What are the chances of drawing 4 aces in a row without replacement in a 52 card deck of cards? What is the probibility of getting six heads when you flip a coin 7times? What is the function of the common bileduct? If you draw a card from a normal deck of cards what is the probability that both cards are the 7 of diamonds? What is the probability of getting a sum greater than 5 if you roll 2 dice? What conditions must be met in order to use the normal distribution to approximate the binomial distribution? How many different ways can you arrange four pictures on a wall if you want them in a straight line horizontally? Which on of the following is not a valid probability a 0.025 b 3 7 c 4 3 d 0.875? What is one advantage of transforming X values into z-scores? When three dice are thrown together what is the probability of getting a total score of 10? What is the Disadvantage of means deviation?